# ExAMPC: the Data-Driven Explainable and Approximate NMPC with Physical Insights

### v1

@@ Abstract
Amidst the surge in the use of Artificial Intelligence (AI) for control purposes, classical and model-based control methods maintain their popularity due to their transparency and deterministic nature. However, advanced controllers like Nonlinear Model Predictive Control (NMPC), despite proven capabilities, face adoption challenges due to their computational complexity and unpredictable closed-loop performance in complex validation systems. This paper introduces ExAMPC, a methodology bridging classical control and explainable AI by augmenting the NMPC with data-driven insights to improve the trustworthiness and reveal the optimization solution and closed-loop performance’s sensitivities to physical variables and system parameters. By employing a low-order spline embedding to reduce the open-loop trajectory dimensionality by over 95%, and integrating it with SHAP and Symbolic Regression from eXplainable AI (XAI) for an approximate NMPC, we enable intuitive physical insights into the NMPC’s optimization routine. The prediction accuracy of the approximate NMPC is enhanced through physics-inspired continuous-time constraints penalties, reducing the predicted continuous trajectory violations by 93%. ExAMPC enables accurate forecasting of the NMPC’s computational requirements with explainable insights on worst-case scenarios. Experimental validation on automated valet parking and autonomous racing with lap-time optimization NMPC, demonstrates the methodology’s practical effectiveness in real-world applications.
@@ I INTRODUCTION
While data-driven control approaches such as Reinforcement Learning (RL) have gained significant traction in academia and robotics due to their minimal system knowledge requirement and their easiness to implement and use, safety-critical applications like Autonomous Driving (AD) demand safe, explainable, and transparent controllers. Although Nonlinear Model Predictive Control (NMPC) inherently offers these trustworthy qualities, it poses implementation and maintenance challenges for non-experts as physical insights are required, particularly when underperformance occurs due to model mismatch or real-time operation failures. A recent survey by [1] explores the combination of RL and MPC to complement their strengths. Additionally, control engineers struggle to calibrate controllers for specific CL Key Performance Indicators (KPIs) that are not directly and analytically linked to the NMPC parameters but rather emerge from the interaction of the controller with the plant, environment and other unmodelled controllers. We propose leveraging operational CL data to: 1) approximate the NMPC using physics-inspired techniques, 2) expose the NMPC’s decision making process within the specific environment conditions, and 3) predict and explain the complex system-level CL performance KPIs.
Training data-hungry learning-based controllers can be impractical and unsafe for systems where only operational data can be collected without disrupting the system. Digital Twins (DTs) enable broader data augmentation and more exploration than the real-world. We propose using a small dataset to elucidate and predict system performance around current operating conditions. This approach accelerates the NMPC design and calibration for the specific operating applications, by integrating Machine Learning (ML)-based prediction and explainability with classical control methods like NMPC, complementing rather than replacing Artificial Intelligence (AI) and NMPC strengths. There exists several approaches that allow transparency and explainabilty of AI models, known as eXplailanable Artificial Intelligence (XAI). Those include Symbolic Regression (SR[2] that provides formulas linking outputs to input features with fast inference. Post-hoc XAI methods like SHAP [3] are also beneficial for interpreting feature contributions in model predictions, particularly for Regression Trees [4].
@@
The paper is organized as follows: in Sec. II we briefly present related work on data-driven and approximate NMPC. In Sec. III we provide a background on the continuous-time optimal control problem, the employed ML regression methods in this work, and on related work combining AI with NMPC. In Sec. IV we introduce the low-order embedding of time-series and the physical-inspired regression model that builds the approximate NMPC for predictions with minimal continuous-time constraints violation and provide an explainability study on the OL prediction of the approximate NMPC. We follow with Sec. V where we introduce the explainable CL performance monitor for the NMPC and demonstrate it on an autonomous driving and racing applications before concluding in Sec. VI.
@@ II Related work
The importance of explainable data-driven control, as discussed in [12], lies in its ability to enhance the transparency of the decision-making process in complex systems. Traditional approaches like Explicit MPC [13], were a first attempt at enabling the real-time execution of the MPC by creating an exact surrogate of the linear MPC through Multi Parametric Programming to precompute the control laws and store them in Look-Up Tables (LUTs). However, this method is often memory-inefficient and unsuitable for nonlinear MPC. While Explicit MPC provides a framework for understanding system behavior in an OL fashion and was extended to provide complexity certification for a particular set of Quadratic Programming (QP) solver as in [14], it falls short in flexibility to systems where the underlying solver in unknown, and lacks the adaptability to forecast performance measures such as computation time of NMPC. We aim at extending the previous approaches into a method allowing nonlinear constraints and dynamics handling, as well as complicated system-level CL KPIs and certification measures as shown in Figure 1.
@@ III-A Continuous-time Optimal Control Problems
An Optimal Control Problem (OCP) is initially posed in continuous time and seeks to optimize a cost function $J(x,u)$ while satisfying a set of (nonlinear) inequality and equality constraints as in the nonlinear continuous Bolza problem [15]:
[…]
@@
where $t\in\mathbb{R}$ denotes time, $x\in\mathbb{R}^{N_{x}}$ is a state of the system, $u\in\mathbb{R}^{N_{u}}$ is a vector of control inputs. The function $\Phi:\mathbb{R}^{N_{x}}\rightarrow\mathbb{R}$ is the terminal cost function and $\mathit{L}:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}$ is the running or stage cost. The continuous-time system dynamics are given by the function $f:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}^{N_{x}}$. The function $g:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}^{N_{g}}$ is the path linear and nonlinear constraints function, and $g_{f}:\mathbb{R}^{N_{x}}\rightarrow\mathbb{R}^{N_{g_{f}}}$ is the terminal constraints function. Finally, $\bar{x}_{0}$ is an input parameter to the OCP setting the initial states condition $x(t_{0})$. The OCP in (1) is solved in a receding horizon fashion resulting in an NMPC CL control framework: at every control iteration of step size $T_{s}$, new state measurements or estimations are fed into the OCP to solve for (1) over a horizon $T_{H}=t_{f}-t_{0}$ and the first control action $u(t_{0})\rightarrow u(t_{0}+T_{s})$ is applied.
@@
We define the orthogonal collocation scheme based on the Trucanted Legendre Series (TLS) of degree $M$ as in (2) to approximate the solutions $x(\tau)$ and $u(\tau)$ of the continuous time OCP in a compact representation through the coefficients $\bm{\alpha}$ rather than a discrete set of points, over the normalized time horizon $\tau\in[-1,1]$ of $t\in[t_{0},t_{f}]$:
$x(\tau)=\sum_{j=0}^{M}\alpha^{x}_{j}\mathcal{L}_{j}(\tau)=(\bm{\alpha}^{x})^{\top}\mathbf{L}_{M}v(\tau).$ (2)
@@
The matrix $\mathbf{L}_{M}\in\mathbb{R}^{(M+1)\times(M+1)}$ is formed by the coefficients of $\mathcal{L}_{j}$ with respect to $\tau$. In particular, the spanning basis are Legendre polynomials $\mathcal{L}_{j}(\tau)$, which have the fundamental orthogonality property. The orthogonality property renders the basis terms independent from each other and less sensitive to perturbation. Moreover, the TLS is parametrized by the coefficients $\bm{\alpha}^{x}=\begin{bmatrix}\alpha^{x}_{0}&\cdots&\alpha^{x}_{M}\end{bmatrix}^{\top}\in\mathbb{R}^{(M+1)\times N_{x}}$ , $\bm{\alpha}^{u}\in\mathbb{R}^{(M+1)\times N_{u}}$, and $v(\tau)=\begin{bmatrix}1&\tau&\tau^{2}&\cdots&\tau^{M}\end{bmatrix}^{\top}$ is a vector with a geometric progression of the normalized time instance $\tau=(2t/t_{f}-1)$ with $t_{0}=0$. Equations (2) and (3) also apply to control trajectories $u(\tau)$ with $\bm{\alpha}^{u}$.
@@
For non-smooth systems and solutions that cannot be fit with a single TLS, the normalized NMPC horizon $[-1,1]$ can be divided into smaller finite elements to create a piecewise polynomial, namely a Legendre-Spline. On each element, every state and input is parametrized by a TLS, the coefficients of which serve as optimization variables $\bm{\alpha}$:
$x(\tau)=\begin{cases}\sum_{j=0}^{M}\alpha^{x}_{1,j}\mathcal{L}_{j}(\tau^{*}),&\tau_{1}=-1\leq\tau<\tau_{2}\\ \vdots\\ \sum_{j=0}^{M}\alpha^{x}_{N_{S},j}\mathcal{L}_{j}(\tau^{*}),&\tau_{N_{S}}\leq\tau\leq 1\end{cases}$ (3)
where $\tau^{*}$ converts the section limits back into $[-1,1]$.
@@
In the generic case we refer to the Legendre-Spline coefficients of a state $x$ as $\alpha^{x}_{i,j}$ where $i\in[1,N_{S}]$ is the element number and $j\in[0,M]$ is the coefficient of order $j$ at this element. A TLS is by design composed out of one single element, and is described by the coefficients $\alpha^{x}_{0,j}=\alpha^{x}_{j}$. The vector containing all the coefficients of section $i$ is denoted as $\bm{\alpha}^{x}_{i}$, and the concatenation of the coefficients over all the sections that represent the Legendre-Spline are denoted as $\bm{\alpha}^{x}$. Furthermore, in [16, Theorem 1], it is proven that the continuous-time trajectory of the state, control trajectories and linear/nonlinear constraints over them can be bounded in a finite approach through the regional convex hulls $\mathcal{P}^{k}$ of the TLS and subsequently the Legendre-Spline through constant matrices $C_{M}^{k}$. For every TLS, the time interval $[-1,1]$ is divided into $K$ regions that are not necessarily equidistant, for less conservatism on the TLS’s extrema approximation:
$\min\{g(\mathcal{P}_{x}^{k},\mathcal{P}_{u}^{k})\}-\epsilon\leq g(x,u)\leq\max\{g(\mathcal{P}_{x}^{k},\mathcal{P}_{u}^{k})\}+\epsilon,$ (4a)

### v2

@@ Abstract
Amidst the surge in the use of Artificial Intelligence (AI) for control purposes, classical and model-based control methods maintain their popularity due to their transparency and deterministic nature. However, advanced controllers like Nonlinear Model Predictive Control (NMPC), despite proven capabilities, face adoption challenges due to their computational complexity and unpredictable closed-loop performance in complex validation systems. This paper introduces ExAMPC, a methodology bridging classical control and explainable AI by augmenting the NMPC with data-driven insights to improve the trustworthiness and reveal the optimization solution and closed-loop performance’s sensitivities to physical variables and system parameters. By employing a low-order spline embedding, we reduce the open-loop trajectory dimensionality by over 95%, and integrate it with SHAP and Symbolic Regression from eXplainable AI (XAI) for an approximate NMPC, enabling intuitive physical insights into the NMPC’s optimization routine. The prediction accuracy of the approximate NMPC is enhanced through physics-inspired continuous-time constraints penalties, reducing the predicted continuous trajectory violations by 93%. ExAMPC also enables accurate forecasting of the NMPC’s computational requirements with explainable insights on worst-case scenarios. Experimental validation on automated valet parking and autonomous racing with lap-time optimization, demonstrates the methodology’s practical effectiveness for potential real-world applications.
@@ I INTRODUCTION
While data-driven control approaches such as Reinforcement Learning have gained significant traction in academia and robotics due to their minimal system knowledge requirement and their easiness to implement and use, safety-critical applications like autonomous driving demand safe, explainable, and transparent controllers. Although Nonlinear Model Predictive Control (NMPC) inherently offers these trustworthy qualities, it poses implementation and maintenance challenges for non-experts as physical insights are required, particularly when under-performance occurs due to model mismatch or real-time operation failures. Additionally, control engineers struggle to calibrate controllers for specific CL Key Performance Indicators (KPIs) that are not directly and analytically linked to the NMPC parameters but rather emerge from the interaction of the controller with the plant, environment and other unmodelled controllers. We propose leveraging operational CL data to: 1) approximate the NMPC using physics-inspired techniques, 2) expose the NMPC’s decision making process within the specific environment conditions, and 3) predict and explain the complex system-level CL performance KPIs.
Training data-hungry learning-based controllers can be impractical and unsafe for systems where only operational data can be collected without disrupting the system. We propose using a small dataset to elucidate and predict system performance around current operating conditions. This approach accelerates the NMPC design and calibration for the specific operating applications, by integrating Machine Learning (ML)-based prediction and explainability with classical control methods like NMPC, complementing rather than replacing Artificial Intelligence (AI) and NMPC strengths. There exists several approaches allowing transparency and explainability of AI models, known as eXplainable Artificial Intelligence (XAI). Those include Symbolic Regression (SR[1] that provides formulas linking outputs to input features with fast inference. Post-hoc XAI methods like Shapley Additive exPlanations (SHAP) [2] are also beneficial for interpreting feature contributions in model predictions, particularly for Regression Trees [3]. By calculating the features’ marginal contributions affecting the prediction outcome of the ML module, SHAP renders the module’s decision making process more transparent and interpretable.
@@
The paper is organized as follows: in Sec. II we briefly present related work on data-driven and approximate NMPC. In Sec. III we provide a background on the continuous-time optimal control problem and the employed ML regression methods in this work. In Sec. IV we introduce the low-order embedding of time-series and the physical-inspired regression model that builds the approximate NMPC for predictions with minimal continuous-time constraints violation and provide an explainability study on the OL prediction of the approximate NMPC. We follow with Sec. V where we introduce the explainable CL performance monitor for the NMPC and demonstrate it on an autonomous driving and racing applications before concluding in Sec. VI.
@@ II Related work
The importance of explainable data-driven control, as discussed in [11], lies in its ability to enhance the transparency of the decision-making process in complex systems. Traditional approaches like Explicit MPC [12], enabled the real-time execution of the linear MPC by precomputing control laws through Multi Parametric Programming and storing them in look-up tables. However, this method remains memory-inefficient and unsuitable for nonlinear MPC. While Explicit MPC provides a framework for understanding system behavior in an OL fashion and was extended to provide complexity certification for a particular set of Quadratic Programming (QP) solver as in [13], it lacks flexibility to systems with unknown solvers, and the adaptability to forecast performance measures such as the NMPC’s computation time. For embedded applications, where the controller’s computation time is critical, [14] propose a method to determine the worst-case execution time of a particular class of linear MPC formulations, allowing an online monitoring and prediction of this KPI. We aim at extending the previous approaches into a method allowing nonlinear constraints and dynamics handling, as well as complicated system-level CL KPIs emerging from the interaction between the controller and the different system components as in Figure 1.
@@
II-A On the use of AI for approximate NMPC
We focus on learning two types of output as in Figure 1:
  1. 1.
    the optimized NMPC OL trajectory for the states’ and control actions’ predicted evolution, which is the OCP (1)’s solution at every instance $t$, for given state estimate $\bm{\mathrm{x}}_{0}(t)$: $\{\bm{\mathrm{x}}_{0}(t),\eta,p(t)\}\rightarrow\{\mathbf{x}_{(\cdot)}(t),\mathbf{u}_{(\cdot)}(t)\}$. The autonomous CL system evolves under uncertainties $\xi$ and scenario parameters $p(t)$ (such as target states, boundary conditions, etc..).
  2. 2.
Existing approaches like [6] that approximate the NMPC solution to predict discrete sequences $\bm{\mathrm{x}}_{(\cdot)}(t)=\{\bm{\mathrm{x}}(t),\bm{\mathrm{x}}(t+T_{s}),\dots,\bm{\mathrm{x}}(t+NT_{s})\}$ over the NMPC horizon length $N$ with a step size $T_{s}$, face several limitations:
  • Data inefficiency: poor scalability with large horizon length and different step sizes when using discrete methods such as direct multiple shooting, despite transformer-based technologies [6, 8] that can handle longer trajectories, but require extensive data.
  • High output dimensionality for discrete sequences.
  • Lack of CTCP, potentially compromising safety.
  • Limited output explainability: individual sequence elements provide minimal insights into trajectory trends, dynamics, and sensitivity to input features and parameters, making them non-intuitive.
The authors of [4] address the first three limitation by embedding the infinite dimensional continuous sequence $\{\bm{\mathrm{x}}_{(\cdot)}(t),\bm{\mathrm{u}}_{(\cdot)}(t)\}$ into a finite set of B-spline coefficients, which have a convex hull property. However, B-splines are complex to construct, highly sensitive to the chosen knot sequence, and the coefficients offer limited physical insights on the trajectory trend and dynamics. Moreover, modeling time-series with high accuracy using B-splines requires a rather dense knot sequence, where poorly chosen knots can lead to ill-conditioning in the fitting problem.
@@ III-A Continuous-time Optimal Control Problems
An Optimal Control Problem (OCP) is initially posed in continuous time and seeks to optimize a cost function $J(\bm{\mathrm{x}},\bm{\mathrm{u}})$ while satisfying a set constraints as in the nonlinear continuous Bolza problem [15]:
$\left\{\begin{aligned} \operatorname*{\textbf{minimize}}_{\bm{\mathrm{x}}_{(\cdot)},\bm{\mathrm{u}}_{(\cdot)}}J(\bm{\mathrm{x}},\bm{\mathrm{u}})&=\Phi(\bm{\mathrm{x}}(t_{f}))+\int_{t_{0}}^{t_{f}}L(\bm{\mathrm{x}}(t),\bm{\mathrm{u}}(t))dt\\ \textrm{subject to }&\dot{\bm{\mathrm{x}}}(t)=f(\bm{\mathrm{x}}(t),\bm{\mathrm{u}}(t)),\\ &g(\bm{\mathrm{x}}(t),\bm{\mathrm{u}}(t))\leq 0,\\ &g_{f}(\bm{\mathrm{x}}(t_{f}))\leq 0,\\ &\bm{\mathrm{x}}(t_{0})=\bar{\bm{\mathrm{x}}}_{0}.\end{aligned}\right$. (1)
@@
where $t\in\mathbb{R}$ denotes time, $\bm{\mathrm{x}}\in\mathbb{R}^{N_{x}}$ is a state of the system, $\bm{\mathrm{u}}\in\mathbb{R}^{N_{u}}$ is a vector of control inputs. The function $\Phi:\mathbb{R}^{N_{x}}\rightarrow\mathbb{R}$ is the terminal cost function and $\mathit{L}:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}$ is the running or stage cost. The continuous-time system dynamics are given by the function $f:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}^{N_{x}}$. The function $g:\mathbb{R}^{N_{x}}\times\mathbb{R}^{N_{u}}\rightarrow\mathbb{R}^{N_{g}}$ is the path linear and nonlinear constraints function, and $g_{f}:\mathbb{R}^{N_{x}}\rightarrow\mathbb{R}^{N_{g_{f}}}$ is the terminal constraints function. Finally, $\bar{\bm{\mathrm{x}}}_{0}$ is an input parameter to the OCP setting the initial states condition $\bm{\mathrm{x}}(t_{0})$. The OCP in (1) is solved in a receding horizon fashion resulting in an NMPC CL control framework: at every control iteration of step size $T_{s}$, new state measurements or estimations are fed into the OCP to solve for (1) over a horizon $T_{H}=t_{f}-t_{0}$ and the first control action $\bm{\mathrm{u}}(t_{0})\rightarrow\bm{\mathrm{u}}(t_{0}+T_{s})$ is applied.
@@
We define the orthogonal collocation scheme based on the Truncated Legendre Series (TLS) of degree $M$ as in (2) to approximate the solutions $\bm{\mathrm{x}}(\tau)$ and $\bm{\mathrm{u}}(\tau)$ of the continuous time OCP in a compact representation through the coefficients $\bm{\alpha}$ rather than a discrete set of points, over the normalized time horizon $\tau\in[-1,1]$ of $t\in[t_{0},t_{f}]$:
$\bm{\mathrm{x}}(\tau)=\sum_{j=0}^{M}\alpha^{x}_{j}\mathcal{L}_{j}(\tau)=(\bm{\alpha}^{x})^{\top}\mathbf{L}_{M}v(\tau).$ (2)
@@
The matrix $\mathbf{L}_{M}\in\mathbb{R}^{(M+1)\times(M+1)}$ is formed by the coefficients of $\mathcal{L}_{j}$ with respect to $\tau$. In particular, the spanning basis are Legendre polynomials $\mathcal{L}_{j}(\tau)$, which have the fundamental orthogonality property. This property renders the basis terms independent from each other and less sensitive to perturbation. Moreover, the TLS is parametrized by the coefficients $\bm{\alpha}^{x}=\begin{bmatrix}\alpha^{x}_{0}&\cdots&\alpha^{x}_{M}\end{bmatrix}^{\top}\in\mathbb{R}^{(M+1)\times N_{x}}$, $\bm{\alpha}^{u}\in\mathbb{R}^{(M+1)\times N_{u}}$, and $v(\tau)=\begin{bmatrix}1&\tau&\tau^{2}&\cdots&\tau^{M}\end{bmatrix}^{\top}$ is a vector with a geometric progression of the normalized time instance $\tau=(2t/t_{f}-1)$ with $t_{0}=0$. Equations (2) and (3) also apply to control trajectories $\bm{\mathrm{u}}(\tau)$ with $\bm{\alpha}^{u}$.
@@
For non-smooth systems and solutions that cannot be fit with a single TLS, the normalized NMPC horizon [1,1][-1,1] can be divided into smaller finite elements to create a piecewise polynomial, namely a Legendre-Spline. On each element, every state and input is parametrized by a TLS, the coefficients of which serve as optimization variables $\bm{\alpha}$:
$\bm{\mathrm{x}}(\tau)=\begin{cases}\sum_{j=0}^{M}\alpha^{x}_{1,j}\mathcal{L}_{j}(\tau^{*}),&\tau_{1}=-1\leq\tau<\tau_{2}\\ \vdots\\ \sum_{j=0}^{M}\alpha^{x}_{N_{S},j}\mathcal{L}_{j}(\tau^{*}),&\tau_{N_{S}}\leq\tau\leq 1\end{cases}$ (3)
where $\tau^{*}$ converts the section limits back into [1,1][-1,1]. The normalized interval is crucial as the Legendre polynomials’ definition and orthogonality properties are derived over it.
@@
In the generic case we refer to the Legendre-Spline coefficients of a state $x$ as $\alpha^{x}_{i,j}$ where $i\in[1,N_{S}]$ is the element number and $j\in[0,M]$ is the coefficient of order $j$ at this element. A TLS is by design composed out of one single element, and is described by the coefficients $\alpha^{x}_{0,j}=\alpha^{x}_{j}$. The vector containing all the coefficients of section $i$ is denoted by $\bm{\alpha}^{x}_{i}$, and the concatenation of the coefficients over all the sections that represent the Legendre-Spline are denoted by $\bm{\alpha}^{x}$. Furthermore, in [16, Theorem 1], it is proven that the continuous-time trajectory of the state, control trajectories and linear/nonlinear constraints over them can be bounded in a finite approach through the regional convex hulls $\mathcal{P}^{k}$ of the Legendre-Spline through constant matrices $C_{M}^{k}$. For every TLS, the time interval [1,1][-1,1] is divided into $K$ regions that are not necessarily equidistant, for less conservatism on the TLS’s extrema approximation:
$\min\{g(\mathcal{P}_{x}^{k},\mathcal{P}_{u}^{k})\}-\epsilon\leq g(\bm{\mathrm{x}},\bm{\mathrm{u}})\leq\max\{g(\mathcal{P}_{x}^{k},\mathcal{P}_{u}^{k})\}+\epsilon,$ (4a)
@@
Within the context of approximate NMPC, we propose the use of RESAFE/COL [16] to overcome the B-spline limitations presented in Sec. II-A (data efficiency, dimensionality, lack of CTCP and explainability) by using spectral collocation with orthogonal basis polynomials to achieve high solution accuracy and impose continuous-time nonlinear constraints by relying on a linear mapping between the coefficients embedding Legendre-Spline and its extrema as in (4). This approach provides interpretable coefficients sequence revealing physical evolution information on the trajectory through the zero-order bias term $\alpha_{0}$ or the $i^{th}$ derivative terms. Moreover, it benefits from a better numerical conditioning and a low-order fitting: as the order $M$ increases, the high order-terms naturally vanish to zero if they do not increase the accuracy, as the coefficients $\alpha_{j}$ decay faster than any polynomial in $j$ [17]. We refer to this method as a low-order embedding.